You put $4,750 into a bank paying 1.4% simple interest per year. If you don’t take anything out of the account, how much do you have after two years?
Answer by mananth(16946)

principal you put = $4750
rate of interest = 1.4 %
time = 2 years
type of interest = simple
Simple interest = p r t / 100
= 4750 * 1.4 * 2 / 100
= 133
Amount at the end of 2 years = principal + interest
= 4750 +133= 4883
You will have $ 4883 at the end of 2 years